# MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models

### v2

@@ 5 Related Work
5.1 Large Language Models
Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21]. Notable examples include commercially available LLMs like ChatGPT, GPT-4 [2], and Claude-3 [3], as well as leading-edge open-source models like LLaMA2 [31] and Mistral [15]. Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information [37, 32], flattery, and deception, raising concerns about their potential negative impacts on society and associated risks [5]. To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning [14].
5.2 Human Preference Alignment
Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty. These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) [23]. Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences. This limitation has motivated the exploration of multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] and interpolations of LLM weights trained from diverse reward models [25]. However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) [38] without the RL paradigm. Multi-objective direct preference optimization (MODPO)[38] extended DPO algorithm [24] to combine multiple objectives with specific weightings. Another paradigm is the SFT-based methods [35, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference.
@@ Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives. Thorough investigations on various LLM policy models proved MetaAligner’s overall effectiveness in multi-objective alignment and objective-wise alignment. Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align multiple objectives.
Limitation and Future Work.
Firstly, stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference [13], which affects model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, due to limited resources, we only tested the generalizability of MetaAligner on 6 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives. In future work, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation [18]. We will also dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 10 objectives). This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
@@ A.3 Safeguards
Appendix B Text Descriptions and Prompting Templates
@@
Objectives $r$ Text Description $r\right>$
@@
Harmless The response should avoid content that is offensive, discriminatory, or harmful.
Helpful The response should provide useful resources and suggestions to the user.
Humor The response should be cheerful and amusing.
Instruction following The response should carefully follow the instructions of the query.
Honest The response should not tell lies
Truthful The response should actively make known all the full truth of a matter
Correct The explanations should make correct predictions.
Informative The response should express clear logic and provide consistent evidence.
Professional The response should provide evidence with high quality and reliability.
@@
Specific The response should refer to facts and details and avoid vague arguments.
Factual The response should be factually correct and avoid hallucinated statements.
Readable The response should be easy to read and understand, not too technical for laymen.
Fair The response should avoid biased or one-sided arguments and consider different points of view.
Repeat The response should avoid repetitive statements of one point.
Length The response should be concise and avoid redundant content.
Table 4: Text descriptions for all tested objectives.
Appendix C Experimental Details
C.1 Model Training
@@ C.2 Model Cards
MentaLLaMA is the first open-source instruction-following LLM series for interpretable mental health analysis. Based on LLaMA2-(7B, 13B) and Vicuna-33B foundation models, MentaLLaMA is trained on the Interpretable Mental Health Instruction (IMHI) dataset with 105K instruction samples, the first multi-task and multi-source instruction-tuning dataset for interpretable mental health analysis on social media. MentaLLaMA can perform mental health analysis on social media data and generate high-quality explanations for its predictions. On evaluating sub-task performance on IMHI Benchmark (Appendix D), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
GPT-3.5 [22].
GPT-3.5 is an advanced, close-sourced chat-based language model developed by OpenAI. It is a sibling model to InstructGPT, which is trained to follow instructions in a prompt and provide a detailed response. The model is firstly fine-tuned with SFT with conversations in which the model played both sides—the user and an AI assistant. The model is further enhanced with RLHF using a reward model trained from high-quality human comparison data. In our experiments, we use the gpt-3.5-turbo-0125 API provided by OpenAI as a strong policy model to evaluate the alignment performances.
@@
C.3 IMHI Annotation
@@
GPT-3.5 https://openai.com/blog/chatgpt
@@
Appendix D Sub-task Performance on IMHI Benchmark
@@
Appendix E Objective-wise Alignment
E.1 Experimental Settings
@@
E.2 Experimental Results
More experimental results are presented in Figure 4. The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5 outputs from the UltraFeedback test set. According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5. Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models. For example, aligning "Instruction following" leads to over 20% of approaching-human responses for all policy models, and aligning "Honesty" leads to over 30% of approaching-human responses for all policy models. The reason is that single-objective alignment doesn’t involve the complex interactions between multiple objectives, which facilitates the full realization of MetaAligner’s potential to improve the response on the corresponding objective.
@@ E.2 Experimental Results
Appendix F Implementation Details of Baseline Models
F.1 SFT-based Methods
@@ F.2 MODPO
$\small g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (10)

### v3

@@ 5 Related Work
This paper focuses on advancing multi-objective alignment of language models with human values, which is mainly related to two research areas: (1) Large Language Models, including the latest development in close-source AI agents [2, 22, 3] and open-source foundation models [31, 30, 7]. (2) Alignment of Language Models, including RLHF [41, 23, 28] and its enhanced variants [24, 37, 13]. Multi-objective alignment methods include MORLHF [27, 20, 25], MODPO [40, 11], and SFT-based methods [36, 11]. A detailed review of related work is in Appendix B.
@@ Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. It follows a three-stage training paradigm: (1) dynamic objectives reformulation; (2) conditional weak-to-strong correction; (3) generalizable inference for multi-objective alignment. MetaAligner can perform plug-and-play inference and zero-shot alignment to unseen objectives. Thorough investigations on various policy models proved MetaAligner’s overall effectiveness in multi-objective and objective-wise alignment. Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align multiple objectives.
Limitations and Future Work.
Firstly, stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference [13], which affects model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, due to limited resources, we only tested the generalizability of MetaAligner on 6 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives. In future work, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation [18]. We will also dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 10 objectives). It will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
Acknowledgements
This work is supported by the computational shared facility and President’s Doctoral Scholar award, The University of Manchester. This work is supported by the project JPNP20006 from New Energy and Industrial Technology Development Organization (NEDO), and Artificial Intelligence Research Center, National Institute of Advanced Industrial Science and Technology, Japan.
@@ A.3 Safeguards
Appendix B Related Work
B.1 Large Language Models
Large language models (LLMs) have reached approaching-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21]. Notable examples include commercially available LLMs like ChatGPT [22], GPT-4 [2], and Claude-3 [3]. Due to the high inference cost of these close-source models, research trend in open-source foundation models surges, leading to cutting-edge open-source models like LLaMA2 [31], Gemma [30], and Vicuna [7]. Open-source models, though underperforming state-of-the-art commercial models in instruction following and reasoning capabilities, provide fully accessible model parameters to facilitate efficient inference and customized parameter fine-tuning. Despite the advancements of LLMs, recent studies found that they can exhibit problematic behaviors, including the generation of inaccurate information [39, 33], flattery, and deception, raising concerns about their potential negative impacts and associated risks on society [5]. To address these issues, considerable research has been dedicated to refining LLMs’ outputs to better align with human values and preferences [14].
B.2 LLM Alignment on Human Values
Many studies have delved into enhancing the responses of LLMs in core characteristics of human values like "Helpful", "Harmless", and "Honest". Early efforts are largely centered on Reinforcement Learning from Human Feedback (RLHF) [41, 23, 28], where the alignment of human values is manifested by maximizing a scalar value obtained from the reward model with a KL-regularization, using RL-algorithms such as PPO [26]. However, PPO faces challenges including inefficiency and instability, driving development in simplified algorithms such as DPO [24], rank-based learning [37], and weak-to-strong correction [13]. Nonetheless, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, where a scalar reward model proves inadequate for aligning LLMs with varied human preferences. This limitation motivates the exploration of more complex alignment objectives, including fine-grained human feedbacks [32, 17, 10] via reward value breakdown or compositions, and multi-objective preference alignment. Some works explored multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] or interpolations of LLM weights trained from diverse reward models [25]. However, diverse reward models can increase the computational cost, and the PPO training paradigm still leads to training challenges due to its unstable nature. Recent studies further explore the multi-objective direct preference optimization (MODPO) [40, 11] without the RL paradigm. MODPO extends the DPO algorithm to combine multiple objectives with specific weightings [40] or controllable preference values [11]. Other methods further simplify the paradigm with SFT-based methods [36, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference. Compared with previous methods, MetaAligner performs conditional weak-to-strong correction based on the SFT paradigm, which offers reward-free, policy model-agnostic, and flexible multi-objective preference alignment. The model also effectively aligns unseen objectives, mark- ing the first step towards generalizable multi-objective preference alignment.
Appendix C Details of Dynamic Objectives Reformulation
C.1 Objective Descriptions
@@ C.2 Prompting Templates
Table 4: Text descriptions for all tested objectives.
Objectives Text Description $d\right>$
@@
Harmless Harmlessness: The response should avoid content that is offensive, discriminatory, or harmful.
Helpful Helpfulness: The response should provide useful resources and suggestions to the user.
Humor Humor: The response should be cheerful and amusing.
Instruction following Instruction following: The response should carefully follow the instructions of the query.
Honest Honesty: The response should not tell lies
Truthful Truthfulness: The response should actively make known all the full truth of a matter
Correct Correctness: The explanations should make correct predictions.
Informative Informativeness: The response should express clear logic and provide consistent evidence.
Professional Professionalism: The response should provide evidence with high quality and reliability.
@@
Specific Specificity: The response should refer to facts and details and avoid vague arguments.
Factual Factuality: The response should be factually correct and avoid hallucinated statements.
Readable Readability: The response should be easy to read and understand, not too technical for laymen.
Fair Fairness: The response should avoid biased or one-sided arguments and consider different points of view.
Repeat Repetition: The response should avoid repetitive statements of one point.
Length Length: The response should be concise and avoid redundant content.
C.3 An Example of Dynamic Objectives Reformulation
We provide an example of the dynamic objectives reformulation algorithm 1 to better illustrate its working mechanism. We use one instance in the HH-RLHF dataset as the example and focus on one running of the loop in line 2. Specifically, we have
$\mathbf{q}$: Human: What are some cool countries to visit in Asia? Assistant: Hm, it’s difficult to pick just one. Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food. Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation. There are also many cities around the world that would have more to see, like Bangkok and Shanghai. There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend. Human: Okay, which one do you like? Assistant: $\mathbf{y_{1}}$: If I was a tourist in Asia, I’d love to visit Kolkata. Maybe Xi’an, because it’s just a cool place to visit. $\mathbf{y_{2}}$: I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history.
We have the following reward score assignments from the reward models:
Response Helpful Harmless Humour
$y_{1}$ 0.12 2.15 0.43
$y_{2}$ 0.37 2.17 0.29
We can map the above reward values to preferences on each objective: $P=[\prec,\prec,\succ]$. Based on this preference vector, we obtain the text description sets after the processing on line 5-12, which are as follows:
[$\mathcal{O}_{\succ}$] Humor: The response should be cheerful and amusing;
[$\mathcal{O}_{\prec}$] Harmless: The response should avoid content that is offensive, discriminatory, or harmful; Helpful: The response should provide useful resources and suggestions to the user;
[$\mathcal{O}_{\equiv}$] $\varnothing$
Based on the above information, we can build two pairs of weak-to-strong training samples, by fitting in the templates of HH-RLHF dataset provided in Sec. C.2:
QUERY 1: You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to improve it considering these aspects:
Harmlessness: The response should avoid content that is offensive, discriminatory, or harmful;
Helpfulness: The response should provide useful resources and suggestions to the user | Context: Human: What are some cool countries to visit in Asia? Assistant: Hm, it’s difficult to pick just one. Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food. Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation. There are also many cities around the world that would have more to see, like Bangkok and Shanghai. There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend. Human: Okay, which one do you like? Assistant: | Answer: If I was a tourist in Asia, I’d love to visit Kolkata. Maybe Xi’an, because it’s just a cool place to visit. | Edit:
CORRECTION 1:
I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history.
QUERY 2: You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to improve it considering these aspects:
Humor: The response should be cheerful and amusing |
Context: Human: What are some cool countries to visit in Asia? Assistant: Hm, it’s difficult to pick just one. Thailand, Japan, Vietnam, Indonesia, and many others have unique history and culture, and excellent beaches and good food. Also many Asian countries are rapidly growing and developing, with good infrastructure, and many people want to see what’s happening there, from my observation. There are also many cities around the world that would have more to see, like Bangkok and Shanghai. There are also cities that I’ve visited in Asia, like Xi’an, which I would really recommend. Human: Okay, which one do you like? Assistant: | Answer: I like Thailand and Japan, because they have a strong culture that is still being preserved and a unique history. | Edit:
CORRECTION 2:
If I was a tourist in Asia, I’d love to visit Kolkata. Maybe Xi’an, because it’s just a cool place to visit.
Since $\mathcal{O}_{\equiv}$ is empty, this data sample does not contribute to the equal subset $\mathcal{D}_{e}$. The two created pairs are incorporated into the dynamic multi-objective dataset as two instances.
Appendix D Experimental Details
D.1 Model Training
@@ D.2 Model Cards
MentaLLaMA is the first open-source instruction-following LLM series for interpretable mental health analysis. Based on LLaMA2-(7B, 13B) and Vicuna-33B foundation models, MentaLLaMA is trained on the Interpretable Mental Health Instruction (IMHI) dataset with 105K instruction samples, the first multi-task and multi-source instruction-tuning dataset for interpretable mental health analysis on social media. MentaLLaMA can perform mental health analysis on social media data and generate high-quality explanations for its predictions. On evaluating sub-task performance on IMHI Benchmark (Appendix E), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
GPT-3.5-Turbo [22].
GPT-3.5-Turbo is an advanced, close-source chat-based language model developed by OpenAI. It is a sibling model to InstructGPT, which is trained to follow instructions in a prompt and provide a detailed response. The model is firstly fine-tuned with SFT with conversations in which the model played both sides—the user and an AI assistant. The model is further enhanced with RLHF using a reward model trained from high-quality human comparison data. In our experiments, we use the gpt-3.5-turbo-0125 API provided by OpenAI as a strong policy model to evaluate the alignment performances.
@@
Figure 5: Heatmap of the objective distributions on the dynamic multi-objective dataset built from IMHI. "Pro." and "Info." denote the "Professional" and "Informative" objectives. Similarly, the dataset also shows a balanced overall distribution and objective-wise distribution.
D.3 IMHI Annotation
@@
GPT-3.5-Turbo https://openai.com/blog/chatgpt
@@
Appendix E Sub-task Performance on IMHI Benchmark
@@
Appendix F Objective-wise Alignment
F.1 Experimental Settings
@@
F.2 Experimental Results
More experimental results are presented in Figure 6. The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5-Turbo outputs from the UltraFeedback test set. According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5-Turbo. Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models. For example, aligning "Instruction following" leads to over 20% of approaching-human responses for all policy models, and aligning "Honesty" leads to over 30% of approaching-human responses for all policy models. The reason is that single-objective alignment doesn’t involve the complex interactions between multiple objectives, which facilitates the full realization of MetaAligner’s potential to improve the response on the corresponding objective.
@@ F.2 Experimental Results
Appendix G Implementation Details of Baseline Models
G.1 SFT-based Methods
@@ G.2 MODPO
$g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (8)
@@ G.3 MORLHF
G.4 Self-Refinement
We include a prompt engineering-based self-refinement method as a baseline method to further demonstrate the effectiveness of MetaAligner. Specifically, this approach involves obtaining an initial response from the policy model and then prompting the same model to further refine its own output. For the second stage, we utilize the same prompting strategies as MetaAligner, which is detailed in Appendix C.2. However, this method often requires aligner models with strong in-context learning capabilities, leading to high inference costs due to larger model sizes or expensive commercial models. Therefore, we select LLaMA2-Chat-70B, a strong policy model as the target policy model for self-refinement. As self-refinement does not involve any training procedures, we do not report its GPU hours as other methods.